OpenStudy (anonymous):

What is the approximate volume of a cylinder with a radius of 3 ft and a height of 4 ft? Use 3.14 to approximate pi. Express your answer in hundredths. ? ft3

OpenStudy (anonymous):

@Science_ALLY @MathHelpp10x

OpenStudy (anonymous):

@Preetha @lordhelix8th @MasterprojectHD

OpenStudy (anonymous):

Formula: V=r^2h So, 3.14 * 9 = 28.26 Then, 28.26 * 4 = 113.04, Rounded = 113.1 So the approximate volume of a cylinder is 113.1. V = 113.1